in python
Write a program that reads the date of birth from the user as 3 integer values (day, month, year)
and then computes and prints their exact age in days. You need to take into account the leap You
may not use any built-in functions related to date and time!
Example:
Someone born on 01/01/2000 is 365 days * 18 years + 5 days (since the years 2000, 2004, 2008, 2012, and 2016 are all leap years) + 352 days (passed in 2019) = 6,927 days old!
Answer #1

All the explanation is in the code comments. Hope this helps!

Please comment in case of a doubt.

Also, take care of indentation by referring to code screenshots.

Code:

import datetime

# function to check for a leap year
def isLeap(year):
  
if(year%400 == 0):
return True
elif(year%100 == 0):
return False
elif(year%4 == 0):
return True
else:
return False

# get today's date
tmp = datetime.date.today()
# store it in an array of 3 elements
now = [tmp.day, tmp.month, tmp.year]

# birthday input from user
birthday = [0] * 3
birthday[0] = int(input('Enter day of birth: '))
birthday[1] = int(input('Enter month of birth: '))
birthday[2] = int(input('Enter year of birth: '))

# days in each month
days_in_each_month = [0, 31, 28, 31, 30, 31, 30, 31, 31, 30, 31, 30, 31]

# total days
days = 0

# loop for days
if(birthday[0] > now[0]):
now[0] = now[0] + days_in_each_month[birthday[1]]
now[1] = now[1] - 1

# calculate days
days = days + now[0] - birthday[0]

# for months
if(birthday[1] > now[1]):
now[2] = now[2] - 1
  
# add days till 12 month of the birthday year
# and then days of months of the present year
tmp = birthday[1] + 1
while(tmp <= 12):
days = days + days_in_each_month[tmp]
tmp = tmp + 1
  
tmp = 1
while(tmp <= now[1]):
days = days + days_in_each_month[tmp]
tmp = tmp + 1
  
else: # normally add days for each month

tmp = birthday[1] + 1
  
# loop for months
while(tmp <= now[1]):
days = days + days_in_each_month[tmp]
tmp = tmp + 1
  
# run loop for years
tmp = birthday[2] + 1
while(tmp < now[2]):
days = days + 365
tmp = tmp + 1
  
# for leap years
# birth year
if(isLeap(birthday[2]) and birthday[1] <= 2):
days = days + 1
# for next year
for i in range(birthday[2] + 1, now[2]):
if(isLeap(i)):
days =days + 1
# current year
if(isLeap(now[2]) and now[1] > 2):
days = days + 1
  
print(days, 'days old!')

Sample run:

Code screenshots: